CRYOGENIC SOLVENT EFFECTS ON THE INFRARED ABSORPTIONS AND BAND WIDTHS OF CF_{3}H \underline{D. L. CEDENO}, J. PENG, D. REYNOLDS, N. MINA-CAMILDE and C. MANZANARES, Department of Chemistry, Baylor University, Waco, TX 76798.